OpenStudy (anonymous):

2x + 7y = 42. FIND Y INTERCEPT PLEASE. I NEED URGENT HELP. Will give medal.

OpenStudy (luigi0210):

The y-intercept is when x=0

OpenStudy (luigi0210):

So solve this: \[\LARGE 2(0)+7y=42\] Should give you an ordered pair \[\LARGE (0, y_{1})\]

OpenStudy (anonymous):

So the y intercept is.....

OpenStudy (luigi0210):

Do the work man, don't be lazy.

OpenStudy (anonymous):

x-intercept To find the x-intercept, plug in y=0 and solve for x -2x%2B7y=14 Start with the given equation. -2x%2B7%280%29=14 Plug in y=0. -2x%2B0=14 Multiply 7 and 0 to get 0. -2x=14 Simplify. x=%2814%29%2F%28-2%29 Divide both sides by -2 to isolate x. x=-7 Reduce. So the x-intercept is . ------------------------------------------ y-intercept To find the y-intercept, plug in x=0 and solve for y -2x%2B7y=14 Start with the given equation. -2%280%29%2B7y=14 Plug in x=0. -0%2B7y=14 Multiply -2 and 0 to get 0. 7y=14 Simplify. y=%2814%29%2F%287%29 Divide both sides by 7 to isolate y. y=2 Reduce.

OpenStudy (anonymous):

Haha, ok.

OpenStudy (luigi0210):

No one's gonna hold your hand your entire life. >_>

OpenStudy (anonymous):

O_O

OpenStudy (anonymous):

y = -2/7

OpenStudy (anonymous):

*gives luigi medal* can u hold my hand? o.O

OpenStudy (anonymous):

lol